OpenStudy (anonymous):

Identify the correct slope and y-intercept of the equation 5x - 3y = -6

OpenStudy (haseeb96):

for y intercept put x =0 in the given equation get the answer 5(0) -3y =-6 -3y=-6 y=-6\-3 y=2 is the y intercept

OpenStudy (anonymous):

slope = 5 over 3; y-intercept = (0, 2) slope = -5; y-intercept = (0, -6) slope = negative 3 over 5; y-intercept = (0, 2) slope = 5 over 3; y-intercept = (0, -6

OpenStudy (haseeb96):

for slope just make the equation in this form y=mx+b

OpenStudy (haseeb96):

-3y= -5x -6 -(3y) = -(5x+6) 3y = 5x +6 y= 5/3 x + 2 here 5/3 is the slope